### v1

# Magnetic Localization for In-body Nano-communication Medical Systems

Abstract
Nano-machines circulating inside the human body, collecting data on tissue conditions, represent a vital part of next-generation medical diagnostic systems. However, for these devices to operate effectively, they need to relay not only their medical measurements but also their positions. This paper introduces a novel localization method for in-body nano-machines based on the magnetic field, leveraging the advantageous magnetic permeability of all human tissues. The entire proposed localization system is described, starting from 1010 magnetometers to be integrated into the nano-machines, to a set of external wires generating the magnetic field. Mathematical equations for the localization algorithm are also provided, assuming the nano-machines do not execute the computations themselves, but transmit their magnetic field measurements together with medical data outside of the body. The whole system is validated with computer simulations that capture the measurement error of the magnetometers, the error induced by the Earth’s magnetic field, and a human body model assuming different possible positions of nano-machines. The results show a very high system accuracy with localization errors even below 1 cm.
Index Terms:
Flow-guided nano-networks, IoT for health, magnetic field, nano-communications, THz communications, wireless localization.I Introduction
Nowadays, modern medicine is becoming more and more biotechnology-oriented. To have precise information about a human health state, numerous tiny medical devices are placed inside of the body with endoscopy techniques, injection of sensors, or even contrast agents for radiography or magnetic resonance imaging. Some of these diagnostic sensors can even traverse human digestive or cardiovascular systems, gathering valuable information relating to different pathologies. But besides measuring intra-body conditions, to be effective for medical diagnostic purposes, these devices should also have communication and localization capabilities.
The communication capability is required to deliver the acquired information to medical systems outside of the body. Until now, numerous concepts have been already proposed that integrate tiny medical sensors within a larger scale Internet of Nano-Bio Things [1, 2, 3, 4, 5]. Assuming the medical sensors are miniature nano-machines, of cubic micrometers in size or even smaller, their communication can be realized according to the so-called nano-communication paradigm, including a few different mechanisms for information transmission [6]. The first of them is based on electromagnetic (EM) communication, like in typical wireless systems, but using very high frequencies, usually terahertz and optical bands, suitable for micro-scale devices and short-range information transfer, and built with innovative materials such as graphene and other two-dimensional nano-materials [7, 8, 9, 10]. The second popular approach, called molecular communication, exploits molecules, e.g., waves of calcium ions, vesicles, or bacteria as information carriers [11, 12, 13]. Other possibilities include very short-distanced but low-delay Forster Resonance Energy Transfer [14, 15, 16] or the ultrasound (for distances up to a few cm) mechanism [17, 18].
The localization capability, on the other hand, is aimed at finding the exact location of a nano-machine at a chosen moment in time. This is crucial to determine where such a machine takes a measurement, which is particularly important if a specific tissue pathology is detected. The nano-machine might be circulating along blood in the cardiovascular system and its location could constantly change. The localization should be possible not only at the moment of injecting the nano-machine into the body but also later, during its operational time.
While wireless localization is a well-known and developed topic, the existing solutions do not fit well with the case of in-body nano-machines. First, the tiny nano-machines operate in strictly limited energy budget conditions and are not capable of performing complex computations for their position estimation. Second, the EM propagation conditions are quite harsh, as the EM signals are very strongly attenuated in body tissues. Moreover, the attenuation differs depending on the tissue type, i.e., blood, fat, skin, or bones. This is no better with other nano-communication mechanisms, i.e., molecular, FRET, or acoustic ones.
Having this in mind, the contribution of this paper is to propose a novel localization approach for in-body nano-machines based on the magnetic field. Like in many localization systems with reference (anchor) points, the reference wires are generating a constant magnetic field. These wires are located in known positions outside of the human body and the emitted magnetic field is constant in time but decreases with the distance from the wires. As the magnetic field penetrates human tissues very well (the relative magnetic permeability is very close to 1 for all human tissues), it can be exploited for determining the distances of a nano-machine from the wires and later the nano-machine position. Moreover, in the proposed solution, the nano-machines do not perform any computations. They just measure the magnetic field, along with other medical measurements. Later, they transmit both the medical and magnetic field measurements through some intermediate devices to a system outside of the body where all the calculations are performed. We present the whole localization system, starting from the 1010 magnetometers to be integrated into the nano-machines, to the set of wires generating the magnetic field. We provide the equations for the localization algorithm in two versions: for three wires and nine (or more) wires. Then, we assess the localization accuracy with computer simulations taking into account the measurement error of the magnetometers, the error induced by the Earth’s magnetic field, and a human body model assuming different possible positions of nano-machines. The results show a very good system accuracy with localization error even below 1 cm.
The rest of the paper is organized as follows. In Sec. II, we review the related localization approaches. In Sec. III, we present the components of the localization system: nano-machines that can be used for medical diagnostic purposes, micro-magnetometers, and a system of wires for generating magnetic field. Then, in Sec. IV, we define the localization algorithm, both its ranging and lateration phases. Later, in Sec. V, we provide a detailed description of computer simulations validating the localization system and including both the magnetic field generated by the system of wires and the Earth’s magnetic field. We simulate the localization process and we calculate its error for different nano-machine positions according to the accepted 3D human body model. Finally, in Sec. VI, we conclude the paper and point out some further research directions.
II Related Work on In-body Nano-node Localization
Performing wireless localization of mobile nano-nodes inside a human body poses several additional challenges compared with the localization in macro-scale wireless networks. First, radio propagation conditions inside the body are quite harsh, with very high signal attenuation, strongly dependent on the type of tissue. For example, for the frequency of 1 THz, which is typically considered for EM nano-nodes, the signal propagation loss at a distance of 1 mm is about 70 dB for fat tissue, 90 dB for skin and even 150 dB for blood [19], which makes classical receive signal strength (RSS) localization approaches hardly applicable. Second, this extremely high signal attenuation limits the communication range of nano-nodes to about 1-2 mm, so a nano-node flowing with blood is frequently isolated from other nodes, unless it approaches another nano-device.
The number of already proposed localization schemes for nano-networks is very limited. Considering nano-nodes flowing with blood in a cardiovascular system, an inertial positioning system has been recently proposed [20]. In this paper, nano-nodes are assumed to get information about their estimated position from anchor nodes attached to the skin, when they flow close by. Later, during their movement, nano-nodes update their positions based on their velocity and rotation sensors. However, this concept is very challenging, as the blood flow is not laminar all the time, but also turbulent, which makes the position updates inaccurate. Moreover, communication between the skin-attached anchors and nano-nodes is rather not possible directly, because of the distance range usually about a centimeter or more. In another paper [21], the time of flight approach was suggested to calculate the two-way distance between anchors and nano-nodes and later to perform localization. This was, though, considered for a grid of static nano-nodes. This concept was later developed in [22], also for static grid-type scenario, where nano-nodes further away from the anchors were localized based on the positions of nano-nodes closer to the anchors. Two other research contributions [23, 24] focus on the hop-counting approach, but this is again feasible only for static nano-nodes arranged in a grid.
Having in mind these discussed limitations of wireless localization, here, we propose to exploit the properties of a static magnetic field. Its main advantage is the invariability in different types of human tissues. The magnetic permeability, which describes how the static magnetic field penetrates different media, is practically the same, up to the sixth decimal place, in all human tissues, see Tab. I. This means that magnetic field strength values can be used for distance estimation independently on the tissues between a magnetic field transmitter and a receiver, e.g., a nano-node.
for different kind of human tissues [25]
| Material | |
| Free Space | 1.00000000 |
| Air | 1.00000040 |
| Water | 0.99999096 |
| Fat | 0.99999221 |
| Bone | 0.99999156 |
| Blood | 0.99999153 |
| Gray Matter | 0.99999103 |
| White Matter | 0.99999120 |
To the best of the authors’ knowledge, this is the first paper where magnetic field localization is proposed for nano-nodes. There are, however, some published works discussing localization with magnetic fields for networks of larger scale. Localization with the magnetic field generated by external magnets is discussed in [26]. The proposed system might be applied inside a body, but the designed mm-size chips transmitting at 480-520 MHz are too large for nano-networks. A system of magnetic coils working as anchors is considered in [27, 28, 29]. Also there, the in-body devices are not suitable for nano-networks, as both the magnetic sensor and the wireless transponder are mm-size. Magnetic field localization is proposed as well for applications with endoscopy capsules being cm-size, assuming a solenoid [30] or a soft-magnet [31] generating the field from inside of a body and some magnetic sensors on the body. In-body magnet rotation can be also tracked for such endoscopy scenarios [32]. Finally, a thorough review of indoor and outdoor magnetic field positioning systems is given in [33].
It is worth mentioning that the magnetic field is likewise considered to control the movement of in-body micro-robots, e.g., in [34]. Knowing also the popularity of magnetic resonance imaging (MRI) techniques, it seems that soon magnetic field may become a force driving a full spectrum of medical applications for human body inspection and diagnostics.
III Magnetic Localization System
Having in mind the main considered application, i.e., in-body nano-networks, here we propose a magnetic field localization system composed of (a) reference anchors outside of a body, and (b) tiny magnetometers mounted on boards of in-body nano-machines. We also assume that information from nano-nodes is transmitted outside of the body through some intermediate devices like a nano-gateway. The whole localization process starts with reference anchors emitting the magnetic field. Then, the nano-nodes measure the magnetic field with their magnetometers. The nano-nodes do not perform any computations. Instead, they transmit these readings, together with other medical measurements, through the intermediate devices to the outside of the body. This approach keeps the nano-nodes simple and avoids wasting their precious energy for localization data processing. The concept is presented in Fig. 1 and the main system components are discussed in the three following subsections.
III-A Reference Anchors
In localization systems, anchors act as sources of reference signals that enable a localized device to calculate the distances from the anchors. In a magnetic field localization system, the anchors should generate a magnetic field. It has been already suggested to use some coils for this purpose, working like electromagnets [28]. However, the mathematical equations of that solution do not have a closed form and they have to be solved numerically with some initial guess values. Instead, here we propose to use a set of metallic wires, each of them supplied with a constant electric current (direct current DC). According to the Biot-Savart law, such a DC wire generates a constant magnetic field decreasing with the distance from the wire axis. Thus, by measuring the magnetic field, a distance from the wire can be determined. It is the basic principle of the localization algorithm proposed here; its mathematical equations are given in Sec. IV. We present two versions of the algorithm, the first one for just three wires, and the second one, more robust and accurate, for nine or more wires.
III-B Nano-machines with Magnetometers
Mobile nano-nodes form a nano-network operating inside the body, e.g., in the cardiovascular system, collecting medical information relating to different pathologies, from tissue damage to cancer biomarkers [35]. Each nano-node, of size about 10 , is a very simple embedded system powered with its own energy source (e.g., a piezoelectric generator) and limited processing, data storage, and communication capabilities [7, 36].
In this paper, we assume the nano-nodes, in addition to medical sensors, are equipped with magnetometers, i.e., instruments to measure the magnetic field. In general, magnetometers are large-scale devices that can be classified into many types like magnetoresistive, spin-valve, superconducting quantum interference devices, or Hall effect-based. For nano-nodes, the magnetometer size should be much smaller than most of them, so here we propose to take advantage of a very recently proposed solution: a graphene Hall effect magnetometer [37]. The active area of this magnetometer is 10 x 10 , which is comparable with state-of-the-art nano-nodes and hopefully could be even reduced with the progress in this technology. This type of device was already tested experimentally, reporting a steady relative measurement error of 1% for the magnetic field up to 120 mT [37].
The magnetic field is a vector parameter and a single Hall effect magnetometer measures just a single component: , , or . Thus, each nano-node is equipped with three such devices, together measuring all three field components.
When performing a medical measurement that requires additional information where the measurement is taken, the magnetometers’ readings are saved together with the medical ones. Later, all this data is transmitted out of the body through the intermediate devices. The nano-node position is calculated after that, in a medical data center or a doctor’s computer, so that the nano-node processor is not burdened with these calculations.
III-C Intermediate Devices
As mentioned in Sec. II, the tiny nano-nodes circulating in a human body have very restricted capabilities, with a communication range limited to about 1-2 mm. Because of that, a nano-node flowing through the same vein where a nano-gateway is installed has only a limited probability of successful transmission to this device [38]. This topic was already further investigated for so-called flow-guided nano-networks in numerous medical applications like diagnosis of artery occlusion [39], bacterial infections, sepsis, heart attacks, or restenosis [40]. The results show that with a sufficient number of nano-nodes in the network, the probability of a successful information transfer out of the body is close to 100%, but it may take even a few hours [39]. A recent work [41] shows that multi-hop communication between the nano-nodes may additionally improve this information transfer. While the detailed analysis of this communication is out of the scope of this paper, we assume that nano-nodes can deliver their gathered medical and magnetic field data to a medical data center through intermediate devices like a nano-router and a body area network.
IV Localization Algorithm
In this section, we explain the proposed localization algorithm. Our proposed solution shares some similarities with the receive signal strength approach, but here the magnetic flux density is measured instead of radio signal strength. The proposed algorithm has two phases: ranging and lateration. While the ranging is the same for different arrangements of DC wires, the lateration phase depends on the number of wires. The basic approach requires only three wires, one for each plane. The more complex algorithm requires at least nine wires, three for each plane.
IV-A Ranging
In the ranging phase, each DC wire is activated in a sequence. At any moment, only one wire is activated. First, nano-sensors measure the magnetic flux density generated by all DC wires, one wire at a time. Since each nano-sensor has three magnetometers, three measurements are taken simultaneously for each wire, corresponding to three orthogonal components of the flux density. These three measurements are then combined into a single vector, whose magnitude is calculated. Finally, the distance from the wire is calculated with the Biot-Savart law:
where
- distance between the nano-node and the wire
- electric current flowing through the wire
- magnitude of measured magnetic flux density
- magnetic permeability of the medium.
IV-B Trilateration with Minimal Number of Wires
In this scenario, the system is considered to have the minimal required number, i.e., three wires. The wires are aligned along the X, Y, and Z axes of the coordinated system (see the wires X, Y, and Z0 in Fig. 2). The nano-nodes are located in the (+x, +y, +z) octant of the coordinate system, so all three coordinates are always positive values. Each distance between a nano-node and a wire, calculated in the ranging phase, creates a cylinder of possible nano-node positions along the wire. For each wire, only two coordinates are involved. Thus, with three wires, a set of simple three equations can be formed, one for each plane:
where
, , - unknown coordinates of the nano-node
, , - distances from the nano-node to the wires X, Y, Z, respectively.
As the coordinates are positive values, they can be calculated from (2) as:
While this approach theoretically can provide the nano-node position, in practice, it is not robust against measurement errors. When a nano-node is located close to one of the , , or planes (see Fig. 2), one of the estimated distances might be much larger than the other ones. Consequently, one of the root values in (3) might be negative. It is then set to zero, to avoid false imaginary parts of the calculated coordinates.
IV-C Multilateration with Many Wires
A more robust algorithm requires a larger number of reference wires. In the following section, we present an approach feasible when we have at least nine wires. Let us consider that all the wires are parallel to one of the axes of the coordinates system and call them , , or wires, according to the axis of their parallelism.
Now, let us consider a set of parallel wires, e.g., wires, like in Fig. 2. Again, the distances calculated in the ranging phase create cylinders of possible nano-node positions along each wire. Thus, localization with Z wires enables to calculation of the (,) coordinates of the nano-node with the following set of equations:
where
- number of Z wires
, - coordinates of the nano-node estimated with Z wires
- distance from the nano-node to the wire
(, ) - wire coordinates.
Considering the first wire is positioned exactly along the axis, its coordinates are =0 and =0. Then, following the approach presented in [42], and assuming we have at least three wires, the first equation in (4) can be subtracted from the rest of the equations, and after reordering the terms, it results in:
This set of equations can be presented in a matrix form:
Denoting the matrices as
Equation (5) can be simply written as:
Then, this matrix equation can be solved using the standard least-squares approach, obtaining the estimate vector containing the and coordinates of the nano-node:
Now, the analogous approach can be performed for and wires. Having at least three wires and three wires, we can follow the algorithm given in (4)-(10) to obtain (,) and (,) coordinates. Each set of wires enables to calculation of two of the nano-node coordinates only. However, assuming the number of , , and wires is respectively equal to , , and , the final nano-node coordinates can be calculated as respective weighted means:
V Simulation Results
To validate the proposed localization algorithms, a computer simulator is written in Python 3 programming language. The simulator reflects the geometric setting of DC wires being the magnetic sources (Sec. III-A) and nano-nodes equipped with magnetometers operating inside of a human body (Sec. III-B). It also takes into account the Earth’s magnetic field which might be a source of measurement errors (Sec. V-A). For each position of the nano-node inside of the body, the simulator calculates the measured magnetic field coming from each DC wire influenced by the Earth’s magnetic field. With these measurements, the position of the nano-node is estimated based on the proposed localization algorithm (Sec. IV). Finally, the localization error is calculated in comparison to the real position of the nano-node. Discussing it in more detail, each simulation follows the subsequent steps:
1.
Geometry definition. The general geometry of the simulation is created: the position of the DC wires and the relative inclination of the Earth’s magnetic field are established. While the Biot-Savart law (1) is defined for infinitely long DC wires, here, for the setup simplicity, we assume the wires are 1-2 m long (see Fig. 5-6), but their length may be increased in practical scenarios.
2.
Nano-nodes positions and inclinations. A voxel model of a 175 cm human male is implemented to emulate the human body. The model is depicted in Fig. 3; it is composed of 17,381 points in total. Each point represents a potential nano-node position for which magnetic field measurements are performed and localization is attempted. At each point, a nano-node is generated with a random orientation, which is repeated 100 times.
3.
Magnetic field of DC wires. For the chosen electric current, the magnetic flux density generated by a DC wire is calculated at each point of the human body, according to the Biot-Savart law (1). Then, the measurement noise is added, which is generated from a uniform distribution of 1% of the computed flux density. The whole step is repeated for each DC wire generating the magnetic field.
4.
Earth’s magnetic field error. A simulated error, resulting from the difference between the real Earth’s magnetic field and its modeled value, is added to each measured magnetic flux density. The Earth’s field model with its accuracy are discussed in a separate Section V-A.
5.
Localization and its error. The proposed localization algorithm is executed and the position of the nano-node is estimated. The Euclidean distance between the real and estimated positions is calculated as the position error. It is averaged over 100 random nano-node orientations at each point of the human body model.
V-A Earth’s Magnetic Field
In static magnetic field measurements, the Earth’s magnetic field (also called geomagnetic field) must be considered. This field is created naturally in the so-called geodynamo process. As a result, Earth is a large magnetic dipole, and its poles are slowly changing their positions in time. The average magnetic flux density at the Earth’s surface is between 25 and 65 .
A couple of models allow for the estimation of the geomagnetic field. In this paper, we use the commonly known World Mathematical Model (WMM) [43]. It was created by the U.S. National Oceanic and Atmospheric Administration’s National Centers for Environmental Information (NOAA/NCEI) and the British Geological Survey (BGS). The current model version was parameterized in 2020 and is supposed to be valid until 2025. The main goal of the WMM is to represent the planet’s magnetic field for all locations of the globe. What is quite important here, it also provides data on how large the model error is compared with the real Earth’s magnetic field. These errors for three orthogonal components (northern, eastern, and vertical ones) of the magnetic flux density are shown in Tab. II for the 2020 and 2025 years.
For the computer simulator, we calculate the average of the model error for 2020 and 2025 (see the third column of Tab. II). Then, for each nano-node, we randomly generate the WMM errors for all three components as uniformly distributed values from the range -error, +error.
| Earth’s field | WMM error | WMM error | Computer simulator |
| component | in 2020 | in 2025 | error |
| X (northern) [nT] | 127 | 135 | 131 |
| Y (eastern) [nT] | 86 | 101 | 94 |
| Z (vertical) [nT] | 146 | 168 | 157 |
V-B Results
Two primary wire arrangements were investigated during simulations: with three and 15 DC wires. The electrical current for all the wires was 100 A. The created magnetic flux density was never higher than 120 mT inside the human body, as it was the maximum value in the measurement range of the used magnetometers.
The scenario with three wires was chosen, as it is the minimal number of wires when the localization is feasible; the respective algorithm is given in Sec. IV-B. On the other hand, with 15 wires, the more advanced algorithm, based on multilateration, can be verified (Sec. IV-C). This wire arrangement creates a cage-like shape where a patient can stand inside (see Fig. 4) for medical investigation. For each simulation and each nano-node, 100 simulations were performed, and the results were averaged. The localization errors are summarized in Tab. III.
In the three wires scenario, the mean position error equals 1.92 cm with a standard deviation of 0.9 cm. The detailed error map is depicted in Fig. 5. The lowest error is in the center of the phantom. The nano-nodes positioned close to the , , or planes may experience large position errors (see the human model left hand and feet in Fig. 5). It is due to the effect described in Sec. IV-B: the ranging errors may result in negative root values when calculating nano-node positions. While it might be corrected in the algorithm, the position error is still high.
The scenario with 15 DC wires produces much better position estimates; the mean error is only 0.74 cm with a standard deviation of 0.18 cm. With the error map (see Fig. 6), it can be seen that again the lowest error is in the center of the human body, but we do not have large errors on feet or hands, as the effect of negative root values does not take place here. While the mean localization accuracy (below 1 cm) might be very attractive for many medical applications, the simpler scenario with just three DC wires might be found more feasible in practice, while being also quite precise.
| Scenario | 3 wires | 15 wires |
| Errors [cm]: | ||
| Maximum X error | 6.81 | 0.850 |
| Maximum Y error | 3.33 | 0.91 |
| Maximum Z error | 6.35 | 0.59 |
| Maximum position error | 6.87 | 1.47 |
| Mean X error | 1.09 | 0.38 |
| Mean Y error | 1.13 | 0.42 |
| Mean Z error | 0.58 | 0.29 |
| Mean position error | 1.92 | 0.74 |
| Std of position error | 0.9 | 0.18 |
VI Conclusions
In this paper, we presented a new magnetic-field-based localization system suitable for in-body nano-devices. In the system, external electric wires with constant current generate a magnetic field that can be measured by tiny magnetometers mounted on board the nano-devices. These measurements are sent out of the body through a body-area network and are used to calculate the nano-device positions. We proposed two versions of the localization algorithm: for three, and nine or more electric wires. We also explained how to take into account Earth’s magnetic field. We performed computer simulations for both algorithm versions showing that even with errors induced by the magnetometers and Earth’s magnetic field, the obtained localization accuracy can be about 1 cm or better. The proposed solution is a necessary step in a variety of potential applications in future medical diagnostic systems.
Future works extending these studies may go in many different directions ranging from specific system applications through nano-device hardware fine-tuning to new networking architectures. One particularly interesting example is substituting the external DC wires, generating the magnetic field, with in-body magnets. Such magnets are already considered for guiding some nano-particles [44]. The advantage of this solution is the fact that no external reference wires are required, as the source of the magnetic field is inside the body. It is, however, still to be investigated if the magnets’ positions could be sufficiently stable and the fields suitably regular and strong for accurate localization. Another related research line is the study of time characteristics of the magnetic fields generated by the external DC wires. As the static field is required for the proposed localization algorithm, it is worth examining how quickly the wires can be switched on and off, as it affects the frequency of localization measurements. These examples of future studies just indicate a wide range of research topics related to magnetic field applications for in-body medical diagnosis.
Acknowledgement
This work was carried out in the framework of COST Action CA20120 INTERACT. It was also supported by the grant CBET-2039189 of US National Science Foundation and by the Polish Ministry of Science and Higher Education with the subvention funds of the Faculty of Computer Science, Electronics and Telecommunications of AGH University.